in a random sample of 452 people 326 voted yes on proposition W. Can we conclude that more than 70% of the population will vote this way? Compute the p-value.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p =0.70

Ha : p >0.70

= x / n = 326/452=0.721

P0 = 0.70

1 - P0 = 1-0.70=0.30

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

=0.721-0.70 / [(0.70*0.30) /452 ]

z= 0.97

P(z > 0.97) = 1 - P(z < 0.97) = 1-0.8340=0.1660 using z table

P-value = 0.1660